# Let’s Influence Algorithms Together: How Millions of Fans Build Collective Understanding of Algorithms and Organize Coordinated Algorithmic Actions

### v1

@@ 4.3.2. Making Algorithmic Manipulation Tutorials as Simple as Possible
There are many instances where the tools available for fan activities differ depending on the device or platform used. For example, an auto-tapper app for sending love taps during live idol streams is only available to Android users, while Apple users must manually tap their screens at intervals. As a result, this task is mostly handled by Android users, as it’s more cumbersome for those with Apple devices. Similarly, a Weibo-based tool named ”SHARE”, which simplifies ”controlling comments” with a keyword filtering function, is also limited to Android devices. This led to Android users taking the lead in managing and evaluating content related to their idols in a long period.
@@
In summary, the key to an effective algorithmic tutorial lies in its simplicity and understandability. The complexity of these tutorials is often tailored to the specific fanbase, with their loyalty and dedication playing a crucial role. As CF40, a fan who has been a core fan in various fan groups, explained,
@@
”Core fans for idols with highly devoted fanbases typically employ more complex but detailed tutorials. In contrast, core fans for actors whose fans are mostly casual supporters tend to opt for simpler tutorials, sacrificing some level of detail to attract more participation in data labor.” (CF40)
@@
Although this observation is not backed by quantitative data, there is a consensus among all the core fans we interviewed: The less manual intervention a tutorial requires, the more likely fans will use it for data manipulation. If the general fan can engage in algorithmic action through a straightforward, automated setup on their phones or computers, their willingness to participate is generally high. Conversely, tasks requiring more cognitive engagement and physical effort are met with lower enthusiasm.
P42 concluded in one sentence, “In all, as core fans, we need to enable the foolish to learn while motivating the lazy to act.”
4.4. ”Our Idol Only Has Us!” The Fan Algorithmic Emotionality of Collective Algorithmic Action (RQ3)
@@
Rational persuasion alone is often insufficient to fully mobilize general fans. What supports large-scale mobilization is a powerful set of emotionality, which is widely recognized and continually reinforced within the fan communities. In their efforts to motivate ordinary fans towards collective algorithmic actions, core fans often employ emotional elements alongside objective explanations of the data’s importance. This strategy evokes emotional sympathy, encouraging fans to participate in collective algorithmic actions. Within the fanbase, this type of rhetoric is called ”Guilt-Tripping Fans.”
”Guilt-Tripping Fans” involves conveying to fans that their idol is facing significant challenges and relies solely on their support, thus creating a sense of urgency and need. CF8, a core fan, describes this approach:
@@
”It’s crucial for fans to understand the gravity of the situation, both rationally and emotionally. Our idols need us, not just in a figurative sense. Particularly in China, fans invest financially to fulfill a sense of obligation rather than merely seeking happiness. Therefore, fostering this sense of responsibility is key to compel fans to engage in data performance.” (CF8)
@@
However, this ”Guilt-Tripping Fans” rhetoric does not work for all general fans. CF37 points out that since the inflammatory rhetoric used by core fans is not mandatory, they cannot persuade every fan to act in the same way. ”Of course, some general fans are dissatisfied with this, thinking, ’As long as I’m happy, I don’t need to worry about how to help my idol,’” CF37 adds,
”But we actually have to accept the existence of such general fans, as they are only a small fraction of the fan base and do not affect the overall algorithmic emotionality within the fandom. This blending of algorithmic labor with emotionality towards loyalty and love for idols has already become an important part of fan culture and won’t be disrupted by the minority who do not recognize it.” (CF37)
Furthermore, CF37 notes that in fandom. there is a comprehensive organizational strategy in place to ensure that most general fans involved in the fan community can perform algorithmic tasks. CF37 states, ”General fans certainly have the option not to participate in these algorithmic actions or in the algorithm manipulation that we, as core fans, ask they to do. However, when it comes to distributing gifts considering idols from the fan club to general fans, such as priority access to idols’ concert tickets, these algorithmic labor activities are considered as key evaluation criteria.” Through this organizational strategy, the fan community disciplines its members, making fans who engage in algorithmic work the mainstream while those who refuse become the non-mainstream in fandom, unrecognized by the fan community and left on its periphery. In other words, these general fans, who do not participate in the fan-driven collective algorithmic actions, may continue to like their idols but find it challenging to integrate into the corresponding fan community.
@@ 5.1. From Individual Folk Theory to Collective Understanding and Action towards Algorithms
Our research contributes to HCI studies by demonstrating how a minority can decode algorithms, develop practical tutorials, and disseminate this algorithmic knowledge to a broader base of general users without specialized expertise. This transferring of folk algorithmic understanding empowers users to collectively manipulate algorithms in their day-to-day practices. This process enhances the study of folk theories of algorithms, expanding it from the dimension of understanding to the dimension of action and from the individual level to the collective level.
@@
This study corroborates the findings of previous algorithm research. Folk theories of algorithms (DeVito, 2021; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022) provide the theoretical foundation for the core fans to decode algorithms. Core fans continuously gather public information from platforms, professional folk understandings of algorithms from other fan groups, and gossip. As previous scholars have noted, these algorithmic insights from other groups form an external knowledge base that further refines their folk understanding of algorithms (Ticona and TSAPATSARIS, 2023; Zhang et al., 2023). Core fans also observe changes in visibility following algorithm manipulation within their own fan community to refine their understanding of algorithms. The external knowledge base, combined with ongoing internal folk experiments, has led to an increasingly clear understanding of algorithms within the fan community.
This validation process of algorithmic understanding for fan community, directly manifested in the improvement or not of data performance, is interpreted by core fans in several aspects: firstly, whether the tutorial reflects a correct understanding of the algorithm; secondly, whether this tutorial effectively mobilizes the enthusiasm of general fans. In the end, through such a process, algorithm tutorials are validated and become core fans’ reliable cognitive understanding of algorithms, as well as the practical dependence of a large number of general fans. This theoretical process can guide subsequent research on folk algorithm understanding, resistance, and manipulation, guiding scholars to consider several aspects of complete algorithm interaction, such as whether users have formed a folk understanding of the algorithm, whether this understanding can be operationalized, whether this manipulation has produced effective or erroneous results, whether users have relied on these effects to improve their algorithmic practice, and whether reliable validation has been incorporated into their algorithmic knowledge system.
We must point out that, as repeatedly emphasized in the results section, core fans cannot directly understand the platform’s algorithmic mechanisms or access any quantitative data. They must continuously observe changes in idol visibility, which falls entirely within the realm of folk theories of algorithm (Eslami et al., 2019). This perspective is purely from the user’s viewpoint, not from that of a professional algorithm engineer. Despite constant iterations and updates over time and across platforms, the algorithm remains a black box for fans. These iterations merely provide tutorials for algorithmic actions that fans can engage with. In our view, developing such a clear understanding is crucial for transitioning from understanding to action in folk algorithmic theories, especially in mobilizing large-scale collective action.
Moreover, the collective algorithmic understanding and actions of fans remain centralized, determined, and disseminated by a small number of core fans. Zhang et al. (Zhang et al., 2023) argue that fans’ algorithmic imaginaries are co-constructed, and most fans form their algorithmic imaginaries not only through direct engagement with algorithms but also by assimilating second-hand knowledge learned from others. Our empirical data further indicate that fans’ algorithmic imaginations originate from core fans, who play a decisive role in spreading these imaginations to countless general fans and persuading them to engage in algorithmic practices. This finding emphasizes the inequality nature of algorithmic imaginaries within group communication, urging future research to focus more on the genesis and dissemination of algorithmic imaginaries within groups, including interactions between different circles within the group, rather than vaguely understanding it as a part of group culture.
@@ 5.2. From Failed or Small-Scale Actions to Successful Large-Scale Computer-Supported Collective Actions
In previous HCI research, the focus has rarely touched on such large-scale, successful computer-supported collective actions. Our study explores a unique case where millions of or even tens of millions of fans mobilize into collective action to manipulate platform algorithms. This action occurs across multiple fan communities, spans major social media platforms, and extends across several countries. Our study validates and extends the effectiveness of the five stages in computer-supported collective action proposed by Shaw et al (Shaw et al., 2014).
First, core fans recognized an urgent and critical need for algorithmic actions to significantly boost their idol’s visibility. They clearly understood the importance of social media data for their idol. They also recognized that this issue was crucial for all general fans, forming the basis for mobilizing millions of general fans. Moreover, all core fans acknowledged the significance of data performance generated by fans. Every core fan we interviewed consistently invested effort in making large-scale algorithmic collective actions a reality, driven by their belief in the reliability of algorithmic labor.
Then, to find effective and appropriate solutions, they had to develop detailed folk algorithmic theories robust enough to guide collective algorithmic actions across different platforms and even in different countries. This step is the most challenging for core fans, as without a solid theoretical foundation to guide their actions, the actions themselves would be impossible. They must continuously conduct grassroots experiments, learn from the experiences of other fan communities, and gather various pieces of information to iteratively refine their folk algorithmic understanding.
When organizing and executing actions, mobilizing large-scale fan participation and ensuring they understand how to act poses a significant challenge for core fans. On one hand, they employ emotional rhetoric to continuously reinforce the importance of participating in collective algorithmic actions among general fans. In a sense, this emotional rhetoric has evolved into a form of emotional coercion, becoming an accepted norm within the fan community. As previous fan studies have revealed, algorithms within fan communities are often emotionalized (Zheng and Xiao, 2023). Our research further explains how algorithmic actions become a shared emotionality among fans. This also fills a gap in the existing theory of computer-supported collective action, highlighting that the emotionality inherent in the collective action itself is a critical component of collective consciousness (Etter and Albu, 2021). On the other hand, core fans must create easy-to-understand, user-friendly, and nearly fully automated algorithm manipulation tutorials to minimize the effort and time required from general fans, thereby maximizing their engagement and contributions. These two aspects together reduce the free-rider effect in fan-driven collective algorithmic actions, allowing general fans to participate in such actions with minimal effort while simultaneously internalizing these actions as intense emotionality, even considering them a key part of their fan identity. These two aspects have also been supported in previous collective actions studies (Olson Jr, 1971; Bonini and Treré, 2024; Treré and Bonini, 2024), where organizers often need to provide strong emotional motivation and clear, actionable guidelines to mobilize large groups effectively. In this study, these points are extended to algorithms: users need to form an emotional connection to the algorithm, such as associating it with a sense of responsibility, and they also need a clear understanding of the actions required to influence the algorithm.
In this whole process, iteration becomes a crucial element. This echoes Shaw et al.’s key factors for the success of computer-supported collective action, namely the need for continuous reflection, evaluation, and improvement of actions (Shaw et al., 2014). Core fans must (1) stay attuned to changes in the platform, as the platform constantly adjusts its algorithms to prevent fans from disrupting its order through algorithmic practices; (2) understand the evolution of computational tools, as these tools are also continuously advancing, requiring core fans to strategically leverage the right tools to support collective action; and (3) comprehend the dynamics of their fan community to fully mobilize their followers. In this context, the core fans’ folk understanding of algorithms and their organization of algorithmic collective actions are not static but are continuously iterated and updated to sustain large-scale computer-supported collective actions.
Our study offers valuable insights for future computer-supported collective actions, particularly those focused on algorithmic interventions. It highlights the necessary conditions for mobilizing the collective action of millions of users. It is worth noting that the case we discussed spans multiple fan communities, targeting algorithmic actions across several countries and platforms, demonstrating the replicability of our proposed theory of collective algorithmic action to some extent. As highlighted in our previous results, these fan communities also learn from one another and draw on each other’s experiences to develop their own collective actions. Therefore, we encourage future HCI scholars to consider the potential generalizability of our theory of computer-supported collective algorithmic action to other groups and actions.

### v2

@@ 6.2. Making Algorithmic Manipulation Tutorials as Simple as Possible
There are many instances where the tools available for fan activities differ depending on the device or platform used. For example, an auto-tapper app for sending love taps during live idol streams is only available to Android users, while Apple users must manually tap their screens at intervals. As a result, this task is mostly handled by Android users, as it is more cumbersome for those with Apple devices. Similarly, a Weibo-based tool named ”SHARE”, which simplifies ”controlling comments” with a keyword filtering function, is also limited to Android devices. This led to Android users taking the lead in managing and evaluating content related to their idols in a long period.
@@
It is worth noting that during the execution of algorithmic practices, fans occasionally employ tools to ease their workload, transforming algorithmic labor into a semi-automated process. Some of these tools are purpose-built apps designed specifically for fan activities, such as Xingfantuan (an app that enables one-click operations for certain data-related tasks on Weibo) and auto-clickers (which allow Android devices to repeatedly tap specific areas on the screen). However, in most cases, fans repurpose existing features to meet their needs and achieve semi-automation. For example, iPhone users might leverage iOS’s built-in AssistiveTouch or gesture controls for semi-automated tasks, while PC users might combine browser auto-refresh features with YouTube’s autoplay function to repeatedly play videos automatically. In essence, semi-automation represents a mindset rather than reliance on any specific tool. It reflects a core fan tendency to adapt and streamline instructional practices into automated workflows, showcasing their strategic and resourceful approach to algorithmic interventions.
@@
The key to an effective algorithmic tutorial lies in its simplicity and understandability. The complexity of these tutorials is often tailored to the specific fanbase, with their loyalty and dedication playing a crucial role. As CF40, a fan who has been a core fan in various fan groups, explained, ”Core fans for idols with highly devoted fanbases typically employ more complex but detailed tutorials. In contrast, core fans for actors whose fans are mostly casual supporters tend to opt for simpler tutorials, sacrificing some level of detail to attract more participation in data labor.” (CF40)
@@
Although this observation is not backed by quantitative data, there is a consensus among all the core fans we interviewed. The less manual intervention a tutorial requires, the more likely fans will use it for data manipulation. If the general fan can engage in algorithmic action through a straightforward, automated setup on their phones or computers, their willingness to participate is generally high. Conversely, tasks requiring more cognitive engagement and physical effort are met with lower enthusiasm.
The figure shows a tutorial created by fans of Xiao Zhan on how to make multiple reservations for his new drama Cang Hai Zhuan on Weibo’s drama rating platform (similar to a Chinese version of IMDb). After making a reservation, the platform notifies users when the drama airs, so the number of reservations reflects public anticipation for the drama. Generally, each user can only make one reservation. Therefore, fans often register multiple accounts to make additional reservations.
After summarizing and designing a streamlined and straightforward operational process, core fans typically compile it into a ”tutorial” to facilitate understanding among fans. These tutorials generally come in three formats: text, images, and videos. Text-based tutorials are primarily used for very simple operations and are often shared informally between general fans for peer-to-peer teaching. Video tutorials, while capable of providing the most detailed instructions, are less convenient to disseminate and can deter some fans due to their length. Image-based tutorials are the most widely used and shared format, typically created by core fans or fan clubs and circulated by general fans. As CF3 noted, ”Image tutorials are the clearest and most straightforward, making them the easiest for most fans to learn from. Only when some fans struggle to understand the image tutorials do we provide them with video tutorials for further guidance. However, most fans can grasp the instructions through image tutorials alone.” CF42 concluded in one sentence, “In all, as core fans, we need to enable the foolish to learn while motivating the lazy to act.”
7. Findings: ”Our Idol Only Has Us!” The Fan Algorithmic Emotionality of Collective Algorithmic Action (RQ3)
@@
Rational persuasion alone is often insufficient to fully mobilize general fans. What supports large-scale mobilization is a powerful set of emotionality, which is widely recognized and continually reinforced within the fan communities. In their efforts to motivate general fans towards collective algorithmic actions, core fans often employ emotional elements alongside objective explanations of the data’s importance. This strategy evokes emotional sympathy, encouraging fans to participate in collective algorithmic actions. Within the fanbase, this type of rhetoric is called ”Guilt-Tripping Fans.”
@@
”Guilt-Tripping Fans” involves conveying to fans that their idol is facing challenges and relies solely on their support, thus creating a sense of urgency and need. CF8, a core fan, describes this approach: ”It is crucial for fans to understand the gravity of the situation, both rationally and emotionally. Our idols need us, not just in a figurative sense. Particularly in China, fans invest financially to fulfill a sense of obligation rather than merely seeking happiness. Therefore, fostering this sense of responsibility is key to compel fans to engage in data performance.” (CF8)
@@
However, this ”Guilt-Tripping Fans” rhetoric does not work for all general fans. CF37 points out that since the inflammatory rhetoric used by core fans is not mandatory, they cannot persuade every fan to act in the same way. ”Of course, some general fans are dissatisfied with this, thinking, ’As long as I’m happy, I don’t need to worry about how to help my idol,’” CF37 adds, ”But we actually have to accept the existence of such general fans, as they are only a small fraction of the fan base and do not affect the overall algorithmic emotionality within the fandom. This blending of algorithmic labor with emotionality towards loyalty and love for idols has already become an important part of fan culture and won’t be disrupted by the minority who do not recognize it.”
Furthermore, CF37 notes that in fandom. there is a comprehensive organizational strategy in place to ensure that most general fans involved in the fan community can perform algorithmic tasks. CF37 states, ”General fans certainly have the option not to participate in these algorithmic actions or in the algorithm manipulation that we, as core fans, ask them to do. However, when it comes to distributing gifts considering idols from the fan club to general fans, such as priority access to idols’ concert tickets, these algorithmic labor activities are considered as key evaluation criteria.” Through this organizational strategy, the fan community disciplines its members, making fans who engage in algorithmic work the mainstream while those who refuse become the non-mainstream in fandom, unrecognized by the fan community and left on its periphery. In other words, these general fans, who do not participate in the fan-driven collective algorithmic actions, may continue to like their idols but find it challenging to integrate into the corresponding fan community.
@@ 8.1. From Individual Folk Theory to Collective Understanding and Action towards Algorithms
Our findings contribute to the HCI literature by emphasizing the shift from individual folk theories to collective algorithmic understanding and action. This research aligns with and extends existing theories on sociotechnical systems (Rogers, 2004, 2012), demonstrating how collective practices enrich our understanding of human-algorithm interaction and foster agency in digital ecosystems. Specifically, we observe how minority groups decode algorithms, create actionable tutorials, and disseminate this knowledge among general users, thereby expanding the scope of folk theories of algorithms from understanding to actionable collective engagement.
@@
This study corroborates the findings of previous critical algorithm studies research. Folk theories of algorithms (DeVito, 2021; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022) provide the theoretical foundation for the core fans to decode algorithms. Core fans continuously gather public information from platforms, professional folk understandings of algorithms from other fan groups, and gossip. As previous scholars have noted, these algorithmic insights from other groups form an external knowledge base that further refines their folk understanding of algorithms (Ticona and TSAPATSARIS, 2023; Zhang et al., 2023). Core fans also observe changes in visibility following algorithm manipulation within their own fan community to refine their understanding of algorithms. The external knowledge base, combined with ongoing internal folk experiments, has led to an increasingly clear understanding of algorithms within the fan community.
Our study extends the theoretical discourse by illustrating how grassroots experimentation within communities contributes to algorithmic literacy. This aligns with the broader HCI agenda of democratizing algorithmic agency, emphasizing the importance of user-driven innovation and shared knowledge in sociotechnical systems (Freeman et al., 2020).
Core fans cannot directly understand the platform’s algorithmic mechanisms or access any quantitative data. Instead, their folk theories emerge from iterative observation and validation, a process that highlights the opacity of platform algorithms. This iterative engagement reflects the ”sensemaking loops” described in HCI (Aselmaa et al., 2017; Urquhart et al., 2024), where users hypothesize, test, and refine their understanding of complex systems. Our findings underscore the importance of fostering algorithmic transparency to support user sensemaking and reduce reliance on speculative practices. These iterations, while constrained by the black-box nature of algorithms, nonetheless provide practical foundations for mobilizing collective action within fan communities.
Furthermore, our analysis reveals the centralization of algorithmic knowledge within fan communities, raising critical questions about power asymmetries in collaborative systems (Jhaver et al., 2021). This underscores the dual nature of core fans (Zheng and Xiao, 2023; Zhang et al., 2023) as both enablers and gatekeepers of collective algorithmic practices, necessitating further HCI research on the implications of such dynamics for inclusivity and knowledge dissemination.
@@ 8.2. From Failed or Small-Scale Actions to Successful Large-Scale Computer-Supported Collective Actions
Our research contributes to the theoretical understanding of large-scale, computer-supported collective actions in HCI. By extending Shaw et al.’s five stages of collective action (Shaw et al., 2014), we offer empirical evidence of how fan communities adapt these stages to engage with algorithmic systems. This demonstrates the relevance and flexibility of these theoretical frameworks in novel digital contexts, advancing HCI’s understanding of algorithm-driven interactions.
Instead of analyzing isolated actions, our study reveals the systemic processes through which core fans mobilize millions of participants across platforms and countries. This contribution lies in showcasing how users shape and influence algorithmic systems at scale. By highlighting how fan communities integrate emotional motivation with practical strategies, our research expands the discourse on the interplay between emotionality and rationality in collective algorithmic practices (Etter and Albu, 2021; Olson Jr, 1971), providing a new lens for studying user-driven digital interventions.
Core fans identified a unifying need for algorithmic action to enhance their idol’s visibility, illustrating how shared intentionality drives large-scale mobilization. This aligns with HCI theories on collaboration and resource alignment (Rozendaal et al., 2019; Halabi et al., 2015), offering a concrete case of how shared goals facilitate effective collective action. Our study about fandom community also addresses ethical challenges, such as coercion and sustainability, contributing to the critical discourse on designing ethical collective practices in algorithmic contexts.
The iterative refinement of algorithmic practices by fan communities highlights adaptability as a key factor in successful collective actions. This finding resonates with HCI principles of iterative design and user feedback loops (Mirabdolah et al., 2023; Paton et al., 2021), emphasizing the need for continuous evaluation and improvement in digital interventions. By linking user adaptability to collective success, our study enriches HCI’s understanding of dynamic, user-driven processes in sociotechnical systems.
Finally, our research underscores the replicability of collective algorithmic actions across diverse contexts. We suggest that future HCI research explore how these practices can inform other domains, such as civic engagement or social movements, thereby expanding the scope of computer-supported cooperative work (Crivellaro et al., 2019). By examining the interaction between collective agency and systemic constraints, our study offers a nuanced perspective on algorithmic interaction and its sociotechnical implications, broadening the horizons of HCI scholarship.
8.3. From Emotional Engagement to Algorithmic Mastery: Rethinking Fan Practices in the Digital Age
Fan studies have traditionally focused on fan identity, creativity, and cultural production (Duffett, 2013; Hills, 2003). Our research expands this scope by highlighting how fans engage in algorithmic interventions to strategically enhance their idols’ visibility. By analyzing these collective practices, we introduce algorithmic systems as a critical site of fan activity, demonstrating how fans adapt their actions to the logic and constraints of digital platforms. This shift broadens the conceptual boundaries of fan studies to encompass the sociotechnical dimensions of fandom in the algorithmic era.
We contribute to fan theory by uncovering the interplay between emotional and rational dimensions of fan-driven algorithmic practices. While emotional investment in idols provides the primary motivation for participation, our findings reveal that fans leverage rational strategies, such as algorithm analysis, tactical voting, and coordinated social media interactions, to maximize their impact. This duality challenges traditional portrayals of fans as purely emotionally driven actors, positioning them instead as strategic agents capable of navigating complex systems. By bridging emotionality and rationality, our study offers a more nuanced understanding of fan behavior in the digital age.
Our research situates fan-driven algorithmic interventions within the broader discourse on fan labor. While prior studies have explored fans’ creative and promotional labor (Stanfill and Condis, 2014), we highlight a new dimension: algorithmic labor. This form of labor involves technical skills, collaborative coordination, and iterative experimentation, reflecting fans’ evolving roles as active participants in shaping algorithmic outputs. Furthermore, we demonstrate how collective fan actions align with and adapt theories of computer-supported cooperative work (CSCW) and collective action (Shaw et al., 2014), offering a theoretical bridge between fan studies and HCI.
8.4. Ethics in Future Platform Design and Governance
The rise of large-scale, computer-supported collective actions highlights critical challenges and opportunities for platform design and governance. These collective actions, exemplified by fan-driven algorithmic interventions, reflect the growing agency of user communities in influencing algorithmic systems. While such practices can generate positive impacts, such as amplifying marginalized voices or fostering digital participation, they also pose ethical and operational concerns for platform integrity, fairness, and user welfare.
Large-scale algorithmic interventions demonstrate how collective user actions can shape platform dynamics, often bypassing intended algorithmic rules. From a governance perspective, platforms must navigate a fine line between fostering user engagement and mitigating manipulation. HCI design theories (Zimmerman and Forlizzi, 2014; Bardzell et al., 2012) emphasize the importance of incorporating diverse user perspectives into platform policies. Platforms could benefit from engaging user communities in the co-creation of governance frameworks, ensuring that collective actions align with ethical and operational goals while maintaining algorithmic integrity.
The ethical implications of collective algorithmic actions warrant closer attention. These actions often rely on emotionally charged mobilization strategies that can lead to coercion, burnout, or exclusion within communities. Additionally, large-scale manipulation of algorithms can overshadow organic content, distort public discourse, and exacerbate inequality in visibility. Addressing these risks requires a dual approach: tools that detect and monitor strategic (mis)behavior help identify manipulation in real-time, while value-sensitive design principles ensure that platforms proactively embed ethical safeguards to prevent exploitative practices. Drawing from HCI theories on value-sensitive design (Borning and Muller, 2012), platforms should integrate transparency mechanisms that clarify how engagement metrics are calculated, reducing misinformation about algorithmic processes. Moreover, participatory design approaches can involve users in shaping algorithmic policies, fostering collective accountability and mitigating harmful strategic behavior at its root.
A recurring issue in our findings is the difficulty users face in understanding algorithmic processes, not only due to their opacity but also because of their inherent complexity. This lack of clarity forces users to rely on speculative practices to influence outcomes. While transparency alone may not fully resolve this issue, platforms could adopt explainable AI approaches (Nazar et al., 2021) to provide clearer insights into how algorithms function while maintaining system security. Additionally, promoting algorithmic literacy through educational tools and feedback systems could empower users to engage with platforms more responsibly and effectively, helping them navigate algorithmic decision-making beyond surface-level explanations.
The iterative and adaptive nature of fan-driven algorithmic actions underscores the need for platforms to design systems that are resilient to evolving user strategies. Drawing from HCI principles of iterative design and user feedback loops (Mirabdolah et al., 2023; Paton et al., 2021), platforms could implement adaptive algorithms that not only detect and respond to manipulation but also learn from user behaviors to improve their fairness and robustness. This dynamic approach can help balance the competing demands of user agency and platform governance.
Finally, it is essential to acknowledge that collective actions are not inherently good or bad; their impact depends on context and intent. Platforms should strive to distinguish between constructive uses of collective agency, such as civic engagement, and harmful manipulation, such as the distortion of public discourse. Building on theories of ethical technology design (Zimmerman and Forlizzi, 2014; Bardzell et al., 2012), platforms can prioritize interventions that amplify collective benefits while mitigating collective harms. For example, proactive moderation policies and inclusive community guidelines could help channel large-scale actions toward positive outcomes.